OpenStudy (anonymous):

From the 12 albums released by a musician, the recording company wishes to release 5 in a boxed set. How many different boxed sets are possible, assuming that the order in which the albums are chosen for a boxed set is irrelevant.

ganeshie8 (ganeshie8):

its a combination problem : picking 5 albums from 12 albums

OpenStudy (anonymous):

well duh, but how do i figure it out.

ganeshie8 (ganeshie8):

use the combuination formula : \(ncr = \frac{n!}{(n-r)!r!}\) \(12c5 = \frac{12!}{(12-5)! 5!}\)

OpenStudy (anonymous):

okay thank you

ganeshie8 (ganeshie8):

you sure u can do the rest ha ?

OpenStudy (anonymous):

yes. i can. thank you :)

ganeshie8 (ganeshie8):

sounds good :)